# Efficient Modelling of Lyman-$\alpha$ opacity fluctuations during late EoR

### v1

@@ 6 Summary & Conclusions
The large scale fluctuations of Ly-$\alpha$ forest spectra observed in high redshift quasars contain crucial information about the physical properties of the IGM. The modelling of these fluctuations becomes challenging due to the effects of various physical quantities at different scales. Ideally, one would require an extremely high dynamic range simulation in order to accurately characterize the fluctuations. However, this becomes computationally intractable given a wide range of uncertain astrophysical parameters. To this end, fast semi numerical approaches can show us an alternate way providing an optimization between speed and accuracy.
@@
In this work, we developed an efficient semi numerical model to estimate the Ly-$\alpha$ opacity fluctuations during the late and post EoR phase. We further compare our model with the available observational data aiming to constrain the model parameters. Below, we summarize the key points of this work.
@@
  • We calibrate our model with respect to high resolution full hydrodynamic Nyx simulation. This essentially modifies the fluctuating GP relation accordingly to be applicable on coarse resolution grid cell (in our case $4\leavevmode\nobreak\ h^{-1}\mathrm{cMpc}$). Our model self consistently includes the relevant ingredients (i.e., density, temperature, and UVB fluctuations) for generating the opacity fluctuations. We checked that the opacity distribution using the calibrated model can provide a reasonable match with the full hydro skewers. The small differences may arise due to the lack of velocity field correlation in our model. However, this effect is unlikely to be strong and can be explored in a separate future project.
@@
  • Next, we demonstrate the capabilities of our model assuming fiducial ionization and thermal histories which are consistent with the available constraints. We found that along with the underlying density field, large scale temperature fluctuations can also play a role in shaping the opacity fluctuations. Further, the photoionization rate/UVB fluctuations act oppositely to the above two. Hence, one needs a very short effective mean free path ($\lambda_{0}$) to produce sufficient opacity fluctuations after overcoming the density and temperature contribution. We also found that the mean photoionization rate ($\langle\Gamma_{12}^{\mathrm{HI}}\rangle$) mainly controls the mean transmission flux by shifting the opacity distribution.
@@
  • We then compare our model predictions with the opacity fluctuations measured by the extended XQR-30 dataset. The efficiency of our model also allowed us to perform a parameter space exploration study by varying $\lambda_{0}$ and $\langle\Gamma_{12}^{\mathrm{HI}}\rangle$. We found that the bestfit model along with the scatter due to realization can reasonably match with the data distribution. This further enabled us to derive constraints on these parameters. The constraints remained consistent within the errorbars with the previous measurements.
@@
Given the prospective wealth of high quality spectroscopic data at high redshifts in the near future, it is crucial to have efficient and reasonably accurate models for gleaning relevant physical information without major computational bottlenecks. To this end, our model can be useful to study Ly-$\alpha$ opacity fluctuations in a large cosmological simulation volume within a tractable computational cost. Hence, this model can complement the full scale hydrodynamic simulations when there is a requirement of uncertain parameter space exploration. With the availability of upcoming large field-of-view telescope surveys, this will be crucial from a theoretical perspective. Furthermore, this can be useful for the synergetic multi-wavelength study of EoR along with other probes like the 21 cm signal from neutral hydrogen distribution. We plan to exploit these capabilities of our model in future projects.
@@ Appendix A Justification for Gaussian likelihood
Here, we show a plot justifying the choice our Gaussian likelihood for the parameter exploration studies. The plot shows the scatter between two random bins due to different realizations of the sightlines along with a few Gaussian contours. The nature of the scatter is reasonably close to being Gaussian which gives us the confidence to use a standard multivariate Gaussian likelihood in this study.

### v2

@@ 6 Summary & conclusions
The large-scale fluctuations of Ly-$\alpha$ forest spectra observed in high-redshift quasars contain crucial information about the physical properties of the IGM. The modelling of these fluctuations becomes challenging due to the effects of various physical quantities at different scales. Ideally, one would require an extremely high dynamic range simulation in order to accurately characterize the fluctuations. However, this becomes computationally intractable given a wide range of uncertain astrophysical parameters. To this end, fast semi-numerical approaches can show us an alternate way to provide an optimization between speed and accuracy.
@@
In this work, we developed an efficient semi-numerical model to estimate the Ly-$\alpha$ opacity fluctuations during the late and post-EoR phase. We further compared our model with the available observational data, aiming to constrain the model parameters. Below, we summarize the key points of this work.
@@
  • We calibrated our model with respect to high resolution full hydrodynamic Nyx simulation. This essentially modified the fluctuating GP relation accordingly to be applicable on coarse resolution grid cell (in our case $4\penalty 10000\ h^{-1}\mathrm{cMpc}$). Our model self-consistently included the relevant ingredients (i.e., density, temperature, and UVB fluctuations) for generating the opacity fluctuations. We checked that the opacity distribution using the calibrated model can provide a reasonable match with the full hydro skewers. The small differences may arise due to the lack of velocity field correlation in our model. However, this effect is unlikely to be strong and can be explored in a separate future project.
@@
  • Next, we demonstrated the capabilities of our model assuming fiducial ionization and thermal histories that are consistent with the available constraints. We found that, along with the underlying density field, large-scale temperature fluctuations can also play a role in shaping the opacity fluctuations. Further, the photoionization rate/UVB fluctuations act oppositely to the above two. Hence, one needs a very short effective mean free path ($\lambda_{0}$) to produce sufficient opacity fluctuations after overcoming the density and temperature contribution. We also found that the mean photoionization rate ($\langle\Gamma_{12}^{\mathrm{HI}}\rangle$) mainly controls the mean transmission flux by shifting the opacity distribution.
@@
  • We then compared our model predictions with the opacity fluctuations measured by the extended XQR-30 dataset. The efficiency of our model also allowed us to perform a parameter space exploration study by varying $\lambda_{0}$ and $\langle\Gamma_{12}^{\mathrm{HI}}\rangle$. We found that the bestfit model, along with the scatter due to realization, can reasonably match the data distribution. This further enabled us to derive constraints on these parameters. The constraints remained consistent within the error bars with the previous measurements.
@@
Given the prospective wealth of high-quality spectroscopic data at high redshifts in the near future, it is crucial to have efficient and reasonably accurate models for gleaning relevant physical information without major computational bottlenecks. To this end, our model can be useful to study Ly-$\alpha$ opacity fluctuations in a large cosmological simulation volume within a tractable computational cost. Hence, this model can complement the full-scale hydrodynamic simulations when there is a requirement for uncertain parameter space exploration. With the availability of upcoming large field-of-view telescope surveys, this will be crucial from a theoretical perspective. Furthermore, this can be useful for the synergetic multi-wavelength study of EoR along with other probes like the 21 cm signal from neutral hydrogen distribution. We plan to exploit these capabilities of our model in future projects.
@@ Appendix A Justification for Gaussian likelihood
Here, we show a plot justifying the choice of our Gaussian likelihood for the parameter exploration studies. The plot shows the scatter between two random bins due to different realizations of the sightlines, along with a few Gaussian contours. The nature of the scatter is reasonably close to being Gaussian, which gives us the confidence to use a standard multivariate Gaussian likelihood in this study.